OpenStudy (anonymous):

The sum of a man's age and his grandmother's age is 102. Eighteen years ago his grandmother was 10 times as old as he was then. What are their present ages?

OpenStudy (ajprincess):

Let the man's present age be x and grandmother's present age be y. x+y=102 y-18=10(x-18) y-18=10x-180 y=10x-162 x+10x-162=102 Can u solve for x?

OpenStudy (anonymous):

So 10x is negative?

OpenStudy (ajprincess):

no.

OpenStudy (ajprincess):

See if u can solve nw. 11x-162=102

OpenStudy (ajprincess):

x is age. so 10x can not be negative. do u get it?

OpenStudy (anonymous):

why do you add it to x?

OpenStudy (ajprincess):

It is given that the sum of present ages of the man and his grandmother is 102. I assumed man's present age to be x nd grandmother's present age to be y. So x+y=102. Is that clear?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

Nw can u tell me the age of the man 18 yrs ago in terms of x and number?

OpenStudy (anonymous):

x-18?

OpenStudy (ajprincess):

yup. vat about grandmother's age 18 yrs ago?

OpenStudy (anonymous):

60?

OpenStudy (ajprincess):

no. I meant in terms of y and number.

OpenStudy (anonymous):

oh, y-18

OpenStudy (ajprincess):

u r right. It is given that the age of grandmother 18 yrs ago was 10 times the age of the man's age 18 yrs ago. So y-18=10(x-18) Is that clear?

OpenStudy (anonymous):

yes. So y-18=10x-180

OpenStudy (ajprincess):

ya.

OpenStudy (anonymous):

So y=10x-162

OpenStudy (ajprincess):

ya

OpenStudy (ajprincess):

Nw subtitute that in x+y=102 and solve for x to find the man's present age. Can u?

OpenStudy (anonymous):

x+10x-162=102 11x-162=102 11x=264 x=24

OpenStudy (ajprincess):

ya perfect. Nw can u find y?

OpenStudy (anonymous):

24-18=6(man's age 18 years ago) 6x10=60(Grandmother's age 18 years ago) 60+18=78(grandmother's present age)

OpenStudy (ajprincess):

perfect.:)

OpenStudy (anonymous):

oh my god. thank you so much for your help!! :)

OpenStudy (ajprincess):

yw.:)